### v2

# When Deepfakes Look Real: Detecting AI-Generated Faces with Unlabeled Data due to Annotation Challenges

Abstract
Existing deepfake detection methods heavily depend on labeled training data. However, as AI-generated content becomes increasingly realistic, even human annotators struggle to distinguish between deepfakes and authentic images. This makes the labeling process both time-consuming and less reliable. Specifically, there is a growing demand for approaches that can effectively utilize large-scale unlabeled data from online social networks. Unlike typical unsupervised learning tasks, where categories are distinct, AI-generated faces closely mimic real image distributions and share strong similarities, causing performance drop in conventional strategies. In this paper, we introduce the Dual-Path Guidance Network (DPGNet), to tackle two key challenges: (1) bridging the domain gap between faces from different generation models, and (2) utilizing unlabeled image samples. The method features two core modules: text-guided cross-domain alignment, which uses learnable prompts to unify visual and textual embeddings into a domain-invariant feature space, and curriculum-driven pseudo label generation, which dynamically exploit more informative unlabeled samples. To prevent catastrophic forgetting, we also facilitate bridging between domains via cross-domain knowledge distillation. Extensive experiments on 11 popular datasets, show that DPGNet outperforms SoTA approaches by 6.3%, highlighting its effectiveness in leveraging unlabeled data to address the annotation challenges posed by the increasing realism of deepfakes. [^1] The code will be open-sourced upon publication.
Introduction
The rapid rise of deepfakes(Zhou et al. 2023; Yu et al. 2023; Zhang et al. 2023, 2024a; Pan et al. 2024; Zhang et al. 2024b; Liu, Ye, and Du 2024), especially in the form of face forgery, has emerged as a major challenge to media authenticity. Face forgery, which involves manipulating or generating human faces in images and videos, poses significant risks across various sectors, including politics, security, and entertainment. These generated media are often so realistic that they are nearly indistinguishable from real footage, eroding public trust in visual content and raising concerns about their potential misuse. As a result, detecting deepfakes, particularly face forgery, has become a critical area of research in artificial intelligence and digital forensics.
Figure 1: Comparison between traditional labeled data training and the new proposed unlabeled data training setting.
Traditional deepfake detection methods(Nguyen et al. 2024; Larue et al. 2023; Qiao et al. 2024; Lin et al. 2024a; Luo et al. 2023) rely heavily on labeled training data, where human annotators manually classify images or videos as either AI-generated or real. While these supervised methods have been effective to some extent, their limitations are becoming more apparent as deepfake generation techniques advance. As AI-generated faces become increasingly realistic, human annotators struggle to make accurate distinctions, rendering the labeling process both time-consuming and less reliable. This reliance on labeled data creates a significant bottleneck, limiting the scalability and practicality of current detection systems, especially as the volume of digital content continues to grow exponentially. Given these challenges, there is an urgent need for deepfake detection methods that can effectively leverage large-scale unlabeled data. Such approaches could alleviate the burden of manual annotation and enable more scalable systems. The vast amounts of unlabeled data available from online social networks offer a unique opportunity to train deepfake detection models without relying on expensive labeled datasets.
Can traditional unsupervised learning methods handle this task? A key challenge in deepfake detection is that faces generated by different AI models closely mimic the distribution of real human faces and are often highly similar to each other. Unlike typical unsupervised learning tasks, where semantic categories are well-separated, faces produced by generative models share many common features with real faces, leading to significant overlap. This overlap complicates the task for traditional unsupervised methods, which rely on clearly defined categories to differentiate between real and fake. As a result, existing unsupervised learning(Bai et al. 2024; Yu, Huang, and Zhang 2025; Zhuang et al. 2022; Deng et al. 2025; Zhang et al. 2025) approaches struggle to capture the subtle differences between real and fake faces, resulting in lower performance and reduced effectiveness in practical applications.
In this work, we introduce the Dual-Path Guidance Network(DPGNet), a novel framework designed to tackle the challenge introduced above. Unlike traditional methods that rely solely on labeled data, DPGNet combines two paths. First, we retain the original labeled data from traditional training settings as the source domain. The second path takes advantage of large-scale unlabeled data, often sourced from online social networks, which reflects the abundant real-world data available for training. DPGNet addresses two main challenges: (1) bridging the gap between labeled source data and the diverse, unlabeled data generated by different AI models, and (2) effectively utilizing large-scale unlabeled images. DPGNet consists of two key components: text-guided cross-domain alignment and curriculum-driven pseudo label generation. The first component uses learnable prompts to align visual and textual information into a shared, domain-independent feature space. This allows the model to better handle different types of deepfake faces while leveraging textual information. The second component mimics human learning by gradually incorporating and learning from more informative unlabeled samples. Through dynamic threshold supervision, it ensures the model focuses on the challenging samples.
We conduct extensive experiments across 11 datasets, including both cross-domain and cross-method evaluations, to evaluate the effectiveness of DPGNet. Our results show that the proposed method outperforms SoTA methods, achieving a significant improvement of 6.3% in detection accuracy. These findings highlight the ability of our method to effectively leverage unlabeled data in real-world scenarios, overcoming the annotation challenges caused by the increasing realism of deepfakes and providing a scalable solution for face forgery detection in the age of AI-generated media. The main contributions are summarized as follows:
Figure 2: Comparison of methods leveraging unlabeled data. CLIP-Source is trained on FF++ (Rossler et al. 2019a). CLIP+PL(x) variants are fine-tuned using UCDDP pseudo-labels with confidence thresholds $\geq 0.9$, $0.8$, and $0.7$. CLIP+DA uses domain alignment with UCDDP, a small subset sampled from the training sets of the five test datasets.
  • We are the first to effectively leverage unlabeled data for face forgery detection, addressing the growing annotation challenges posed by increasingly realistic AI-generated content and advanced generation techniques.
  • We introduce the DPGNet, a novel framework that combines two paths: one for bridging the gap between labeled source data and diverse, unlabeled data generated by different methods, and another for effectively utilizing large-scale unlabeled data to improve performance.
  • Comprehensive experiments have validated the effectiveness of DPGNet across 11 popular datasets under various settings, demonstrating superior performance (+6.3%).
Figure 3: Framework overview of the proposed DPGNet, illustrating the overall architecture and the interaction between its two core modules: text-guided cross-domain alignment and curriculum-driven pseudo label generation.
Related Work
Our work focuses on leveraging unlabeled data in face forgery detection to bridge the domain gap between source and target domains, preserving prior knowledge while enhancing robustness to various forgery types. We review related works on generalizable deepfake image detection and domain adaptation, highlight their limitations and elaborate on our contributions.
Generalizable deep fake image detection. The emergence of new fake methods has become an important challenge that plagues detector research. To solve this problem, generalization has become the mainstream direction of current research. Some works design detectors by mining general fake clues that may exist in fake images. (Yan et al. 2023; Fu et al. 2025a; Huang et al. 2023; Dong et al. 2023) extract general fake features for learning by decoupling and reconstructing images or IDs. (Li et al. 2021; Luo et al. 2021; Liu et al. 2021a) trains detectors by mining the differences between fake and real images in the frequency domain. (Yan et al. 2024a) uses distillation learning between teacher and student encoders to enable student encoders to learn more general fake knowledge. Some works pursue generalization by simulating more fake methods through operations such as data augmentation/mixing. In recent studies, (Yan et al. 2024a; Yermakov, Cech, and Matas 2025; Cui et al. 2024) leverage the powerful feature extraction capabilities of large visual language models such as clip to achieve efficient generated image detection. (Cui et al. 2024) use adapters to guide clip to pay attention to fake clues. (Yan et al. 2024a) and (Yermakov, Cech, and Matas 2025) respectively proposed two different fine-tuning methods to unlock the potential of clips by exploring their weighted contextual relationships.
Unsupervised Domain Adaptation (UDA). Unsupervised domain adaptation aims to address the domain gap in detection by leveraging unlabeled target domain data. Early UDA methods (Long et al. 2015, 2017; Sun and Saenko 2016) focus on learning domain-invariant features by minimizing domain differences. For example, DAN (Long et al. 2015) uses the maximum mean difference to align domains, while CORAL (Sun and Saenko 2016) employs linear projection to match second-order statistics. Adversarial methods, such as DANN (Ganin and Lempitsky 2015) and CDAN (Long et al. 2018), use domain discriminators to align source and target feature distributions. However, these methods perform poorly in deep face forgery detection due to a key difference: unlike traditional classification tasks with clear semantic boundaries (e.g., ‘cat’ vs. ‘dog’), real and fake faces share the same high-level attributes (‘face’). The difference between them lies in the traces of artifacts (e.g., artifacts in the eye region or inconsistent skin texture) rather than different semantic categories. Pseudo-labeling strategies (Zhou et al. 2024)alleviate this by assigning labels to unlabeled data, but high-confidence pseudo-labels often prioritize simple samples with low value, leading to error propagation and ignoring challenging cases that are critical for robust detection. Some parameter efficient fine-tuning (PEFT) methods (Han et al. 2024), such as LoRA(Hu et al. 2022), achieve domain adaptation by fine-tuning the visual base model, but there is a risk of distorting the pre-trained knowledge (Cai et al. 2019; Tang, Chen, and Jia 2020), which may lead to low ranking (Yan et al. 2024b) in the feature space. Our method overcomes these limitations by integrating visual-language alignment and dynamic pseudo-labeling, effectively capturing various artifact patterns while preserving prior knowledge and ensuring robust generalization across domains.
Method
Problem Definition
The task involves a labeled source domain dataset $\mathcal{D}_{s}=\{(x_{i}^{s},y_{i}^{s})\}_{i=1}^{N_{s}}$ and an unlabeled target domain dataset $\mathcal{D}_{t}=\{x_{i}^{t}\}_{i=1}^{N_{t}}$ , where $x_{s}^{i}$ is an image and $y_{s}^{i}\in\{0,1\}$ represents the label. Our goal is to solve the problem of covariate shift between $\mathcal{D}_{s}$ and $\mathcal{D}_{u}$, and effectively utilize a small number of target domain samples $\mathcal{D}_{t}$ extracted from multiple cross-domain datasets with different distributions, so that the model can generalize well to the full target domain data $\mathcal{D}_{u}=\{x_{t}^{i}\}_{i=1}^{N_{u}}$, where $N_{t}$ and $N_{u}$ represent the number of samples, and through our setting $N_{t}\ll N_{u}$.
To emulate the diverse forgery techniques encountered in real-world scenarios, we constructed two composite datasets, UCDDP and UDF40, by sampling a small subset of images from the training sets of multiple cross-domain datasets $\{\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{K}\}$. This approach ensures diversity in unknown forgery methods and data distributions. Specifically, UCDDP encompasses images generated by various unknown forgery techniques, exhibiting significant distributional variations across samples. In contrast, UDF40 includes forgery methods distinct from those in the source domain while maintaining a consistent data distribution with the source domain. The task is formalized as minimizing the expected classification error on the target domain:
$\min_{f}\mathbb{E}_{\mathbf{x}_{u}\in\mathcal{D}_{u}}\left[\ell(f(x_{u}),y_{u})\right]$ (1)
where $\ell$ denotes the classification loss function and $y_{u}$ represents the latent true label of the sample.
Framework Overview
We conduct a pre-experiment to illustrate the motivation behind our framework. As shown in Figure 2, we evaluate four approaches for leveraging unlabeled data: (1) a baseline excluding unlabeled data, (2) high-confidence pseudo-labeling, (3) domain alignment, and (4) our proposed method. The results reveal that conventional pseudo-labeling often favors easily classified samples, which are typically simplistic pseudo-samples with limited generalization. Additionally, the domain gap causes visual embeddings in $\mathcal{D}_{t}$ to diverge from those of the source domain’s trained model, leading to unreliable feature alignment.
Inspired by this, we propose using text clues as a bridge to coordinate source-target domain knowledge, improve distribution shift, and introduce a curriculum learning strategy to dynamically integrate high-value difficult samples. The DPGNet consists of two stages: source domain pre-training and joint domain adaptation. In the first stage, to establish text-guided domain alignment, given the pre-processed source domain image $x_{s}\in\mathcal{D}_{s}$, we use the visual encoder $E_{v}$ to extract semantically rich features $z_{s}^{f}\in\mathbb{R}^{256\times 1024}$ from it, and align $z_{s}^{f}$ with the real/fake specific text vectors $e_{t}^{f}\in\mathbb{R}^{768}$ generated by the hint learning module. This alignment is optimized through a composite constraint $\mathcal{L}_{\text{Source}}$, ensuring that $z_{s}^{f}$ captures robust, category-independent semantic representations, while $f_{\text{real}},f_{\text{fake}}$ encodes domain-invariant features of real and fake samples.
In the second phase, for unlabeled images from the target domain $x_{t}\in\mathcal{D}_{t}$, the DPGNet extracts features using a fine-tuned encoder $E_{v}$ to obtain $z_{t}^{f}$, perform classification inference, and pad them according to a high confidence threshold $\lambda_{tf}$, thus obtaining a feature base for the target domain $B_{\text{real}},B_{\text{fake}}$. To generate reliable pseudo labels, we measure the feature distances between $z_{t}^{f}$ and $B_{\text{real}},B_{\text{fake}}$ in the feature library and combine them with the classifier predictions to obtain pseudo labels. To improve the quality of pseudo labels, we introduce a curriculum learning strategy to dynamically adjust the screening threshold $\lambda_{lt}$ to merge challenging high-value samples. To mitigate catastrophic forgetting, we apply cross-domain augmentation to source domain features $z_{s}^{f}$ and employ knowledge distillation to align representations of $\mathcal{D}_{s}$ and $\mathcal{D}_{u}$, enabling joint training and enhancing robust generalization across domains.
Text-Guided Cross-Domain Alignment
In this module, the visual encoder is jointly trained with learnable text prompts. Drawing inspiration from prior work (Yermakov, Cech, and Matas 2025), we selectively fine-tune the layer normalization parameters of the first 24 Transformer layers to preserve pre-trained knowledge. For a source domain image $x_{s}\in\mathcal{D}_{s}$, the visual encoder $E_{v}$ extracts visual embedding features $z_{s}^{f}\in\mathbb{R}^{256\times 1024}$, which are processed by a classification head $h(\cdot)$ to produce the prediction $\hat{y}_{s}=h(z_{s}^{f})$. This process enables $E_{v}$ to effectively capture forgery-related features, which can be formulated as:
$\mathcal{L}_{\text{cls}}=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}w_{i}\cdot\text{CE}(\hat{y}_{s,i},y_{s,i})$ (2)
where $w_{i}$ is initially set to 1. To mitigate the imbalance in the sample sources domain and enhance the semantic representation of the real faces, we assign a higher learning weight $w_{i}=2.0$ to the real samples ($y_{s,i}=0$).
Learnable Text Prompts. We introduce two trainable text prompts, initialized as ‘real face photo’ and ‘deep fake face photo’, parameterized as $v_{\text{real}}\in\mathbb{R}^{d}$ and $v_{\text{fake}}\in\mathbb{R}^{d}$, where $d$ is the text embedding dimension. These prompts are fed into the CLIP text encoder $E_{t}$ to generate concept embeddings:
$f_{\text{real}}^{f}=f_{t}(v_{\text{real}}),\quad f_{\text{fake}}^{f}=f_{t}(v_{\text{fake}})$ (3)
where these embeddings serve as semantic anchors for real and fake categories, capturing domain-invariant, independent concepts that transcend source-target distribution shifts.
Visual-Text Alignment. To ensure that the semantics encoded by the visual feature $z_{i}^{f}$ are compatible with $e_{y_{i}}^{f}$, we perform visual-text alignment:
$\mathcal{L}_{\text{alig}}=-\log\frac{\exp(\text{sim}(z_{i}^{f},f_{y_{i}})/\tau)}{\exp(\text{sim}(z_{i}^{f},f_{\text{real}})/\tau)+\exp(\text{sim}(z_{i}^{f},f_{\text{fake}})/\tau)}$ (4)
where $\text{sim}(\cdot,\cdot)$ denotes cosine similarity, and $f_{y_{i}}$ is the text embedding corresponding to the ground-truth label $y_{i}$. This alignment minimizes domain-specific biases, ensuring that the visual encoder’s learned representation $\mathcal{F}_{v}=\{z_{i}^{f}\mid x_{i}\in\mathcal{D}_{i}\}$ is invariant to domain-specific artifacts $\mathcal{A}_{d}\subseteq\mathcal{F}_{v}$ irrelevant to the classification task.
Textual Contrastive Enhancement. To enhance the discriminative power of $\mathcal{F}_{v}$ and promote class cohesion and separation, we apply a contrastive constraints:
$\mathcal{L}_{\text{con}}=\mathbb{E}_{x_{i}\sim\mathcal{D}_{i}}\left[-\text{sim}(z_{i}^{f},f_{y_{i}})+\text{sim}(z_{i}^{f},f_{y_{i}})\right]$ (5)
where $f_{y_{i}}$ is the text embedding of the opposite class. This process prioritizes task-relevant semantic features $\mathcal{S}_{c}\subseteq\mathcal{F}_{v}$ (related to authenticity) over domain-specific features $\mathcal{A}_{d}$.
Curriculum-Driven Pseudo Label Generation
In this module, we design a curriculum learning strategy with dynamic threshold supervision to address the limitations of common pseudo-label-based strategies.
Feature library construction. We extract visual features $z_{t}^{f}=f_{v}(x_{t})$ for all target samples and generate initial pseudo-labels and their confidence scores for each sample through the classification head $h(\cdot)$:
$(\lambda_{\text{real}}^{i},\lambda_{\text{fake}}^{i})=h(z_{t}^{f}),\quad\lambda_{i}=\max(\lambda_{\text{real}}^{i},\lambda_{\text{fake}}^{i})$ (6)
where we retain samples of $\lambda_{i}\geq\lambda_{tf}$ to construct the feature library $\mathcal{B}$. $\mathcal{B}$ is split into real/fake sub-libraries $\mathcal{B}_{\text{real}}$ and $\mathcal{B}_{\text{fake}}$ according to the pseudo labels, where $\mathcal{B}_{\text{real}}$ and $\mathcal{B}_{\text{fake}}$ contain features labeled as real and fake according to the initial visual encoder predictions, respectively, which we use as a ‘simple’ reference case for curriculum learning.
Dynamic Threshold for Pseudo Label Generation. For each target sample $x_{t}^{j}$, we generate pseudo labels through a dual-verification process. First, consistent with the calculation in the feature library construction, we get the CLIP-based prediction $\hat{y}_{\text{clip}}^{j}=\lambda_{i}$. Next, we assess the feature library distance by calculating the minimum L2 distance between $z_{t}^{f}$ and the features in the real and fake sub-libraries:
$\quad d_{\text{fake}}^{j}=\min_{\mathbf{z}\in\mathcal{B}_{\text{fake}}}\|\mathbf{z}-z_{t}^{f}\|_{2}$ (7)
where we assign $\hat{y}_{\text{bank}}^{j}=0$ if $d_{\text{fake}}^{j}>0.5$, or $\hat{y}_{\text{bank}}^{j}=1$ otherwise. A pseudo label $\hat{y}_{unlab}^{j}$ is accepted if $\hat{y}_{\text{clip}}^{j}=\hat{y}_{\text{bank}}^{j}$ and $\hat{y}_{\text{clip}}^{j}\leq\lambda_{lt}^{(t)}$, where $\lambda_{lt}^{(t)}$ is a dynamic threshold. Based on the analysis of simple samples in the feature pool, we initialize $\lambda_{lt}^{(0)}=0.85$, which gradually decreases to 0.70 during training. This curriculum strategy initially prioritizes easier samples and progressively incorporates more challenging samples as $\lambda_{lt}^{(t)}$ decreases, ensuring robust learning from diverse target features $\mathcal{F}_{t}=\{z_{t}^{f}\mid x_{t}\in\mathcal{D}_{t}\}$ while minimizing bias toward less informative samples.
End-to-End Training Strategy
Latent Space Domain Augmentation. Grounded in domain adaptation theory, the transition from source to target domain training often introduces conflicts due to distributional disparities, leading to degraded performance. Therefore, we propose a cross-domain augmentation strategy that integrates latent representations into the source domain’s forgery feature space. By augmenting the source domain’s feature space with target domain information, we expand the latent feature space of training samples and create an intermediate representation that bridges the two domains. This facilitates a smoother learning process, avoiding abrupt shifts between domains. Specifically, we compute a linear combination of latent features $z_{s}^{f}$ and $z_{t}^{f}$, extracted from source samples $x_{s}\in\mathcal{D}_{s}$ and target samples $x_{t}\in\mathcal{D}_{t}$:
$z_{d}^{f}=\alpha z_{s}^{f}+(1-\alpha)z_{t}^{f},\quad i\neq k,\quad\alpha\sim\text{Uniform}(0,1)$ (8)
where $\alpha$ controls the interpolation weights. Learning the augmented features $z_{d}^{f}$ strengthens the decision boundary and preserves shared feature structures across domains. We define this process as follows:
$\mathcal{L}_{\text{dis}}=\mathbb{E}_{x_{s}\in\mathcal{D}_{s},x_{t}\in\mathcal{D}_{t}}\left[\|z_{s}^{f}-z_{d}^{f}\|_{2}^{2}\right]$ (9)
End-to-End Loss Design. The overall training goal integrates the learning of the source domain and the target domain. The first stage is source domain training, and the second stage is joint training. The whole process is end-to-end, formulated as follows:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{p1}}+\mathcal{L}_{\text{p2}}$ (10)
$\mathcal{L}_{\text{p1}}=\mathcal{L}_{\text{cls}}^{\text{s}}+\lambda\mathcal{L}_{\text{alig}}^{\text{s}}$ (11)
$\mathcal{L}_{\text{p2}}=\mathcal{L}_{\text{pse}}+\lambda_{1}\mathcal{L}_{\text{cls}}^{\text{s}}+\lambda_{2}\mathcal{L}_{\text{alig}}^{\text{s}}+\beta\mathcal{L}_{\text{dis}}$ (12)
$\mathcal{L}_{\text{pse}}=\mathcal{L}_{\text{con}}^{\text{t}}+\mathcal{L}_{\text{cls}}^{\text{t}}+\mathcal{L}_{\text{alig}}^{\text{t}}$ (13)
where $\lambda$, $\lambda_{1}$, $\lambda_{2}$ and $\beta$ are weight factors for balancing. The specific settings are detailed in the experimental section.
Experiments
Settings
Datasets.
We used the following nine datasets: FaceForensics++ (FF++)(Rossler et al. 2019a), Deepfake Detection Challenge(DFDC)(Dolhansky et al. 2020a), preview version of DFDC(DFDCP)(Dolhansky et al. 2019), two versions of CelebDF (CDF-v1, CDF-v2)(Li et al. 2020c, b), DeepfakeDetection (DFD)(Deepfakedetection 2021), DF40(Yan et al. 2024c), UCDDP and UDF40, respectively; UCDDP is an unlabeled dataset obtained by sampling a small amount from the training sets of Deepfake Detection Challenge(DFDC), preview version of DFDC (DFDCP), two versions of CelebDF (CDF-v1, CDF-v2), DeepfakeDetection (DFD), and UDF40 is an unlabeled dataset obtained by sampling a small amount from the data subsets of different counterfeiting methods of DF40.
Dataset
Real
Videos
Fake
Videos
Total
Videos
Synthesis
Methods
Total
Image
FF++ (Rossler et al. 2019b) 1000 4000 5000 4 320k
CelebDF-v1 (Li et al. 2020b) $\dagger$ 408 795 1203 1 77k
CelebDF-v2 (Li et al. 2020b) $\dagger$ 590 5639 6229 1 399k
DFDCP (Dolhansky et al. 2019) $\dagger$ 1131 4119 5250 2 336k
DFDC (Dolhansky et al. 2020b) $\dagger$ 23654 104500 128154 8 8202k
DFD (Deepfakedetection 2021) $\dagger$ 363 3000 3363 5 215k
DF40 (Yan et al. 2024c) * $\sim$1500 0.1M+ 0.1M+ 40 1M+
UCDDP $\dagger$ 10+ 18k
UDF40 * 6 9k
Table 1: Details of the dataset used. UCDDP and UDF40 are uniformly sampled from the above datasets. The symbols * and $\dagger$ represent the corresponding sampling relationship.
Methods Venue Backbone CDF-v1 CDF-v2 DFD DFDC DFDCP Avg.
EfficientB4 (Tan and Le 2019) PMLR’19 EfficientNet 0.791 0.749 0.815 0.696 0.728 0.756
Xception(Chollet 2017) ICCV’19 Xception 0.779 0.737 0.816 0.708 0.737 0.755
Face X-ray (Li et al. 2020a) CVPR’20 HRNet 0.709 0.679 0.766 0.633 0.694 0.696
F3Net (Qian et al. 2020) AAAI’20 Xception 0.777 0.735 0.798 0.702 0.735 0.749
FFD (Dang et al. 2020) CVPR’20 Xception 0.784 0.744 0.802 0.703 0.743 0.755
SRM (Luo et al. 2021) CVPR’21 Xception 0.793 0.755 0.812 0.700 0.741 0.760
SPSL (Liu et al. 2021b) CVPR’21 Xception 0.815 0.765 0.812 0.704 0.741 0.767
Recce (Cao et al. 2022) CVPR’22 Designed 0.768 0.732 0.812 0.713 0.734 0.752
CORE (Ni et al. 2022) CVPR’22 Xception 0.780 0.743 0.802 0.705 0.734 0.753
UCF (Yan et al. 2023) ICCV’23 Xception 0.779 0.753 0.807 0.719 0.759 0.763
ED (Ba et al. 2024) AAAI’24 ResNet-34 0.818 0.864 - 0.721 0.851 -
LSDA (Yan et al. 2024a) CVPR’24 EfficientNet-B4 0.867 0.830 0.880 0.736 0.815 0.826
ProDet (Cheng et al. 2024) NeurIPS’24 EfficientNet-B4 0.909 0.844 - 0.811 0.724 0.822
UDD (Fu et al. 2025b) AAAI’25 ViT-B/16 - 0.869 0.910 0.758 0.856 -
Effort (Yan et al. 2024b) ICML’25 CLIP (ViT-L/14) 0.926 0.878 0.922 0.822 0.835 0.877
ForensicsAdapter (Cui et al. 2024) CVPR’25 CLIP (ViT-L/14) 0.914 0.900 0.933 0.843 0.890 0.896
DPGNet (ours) - CLIP (ViT-L/14) 0.973 0.957 0.951 0.892 0.917 0.938
(↑4.7%) (↑5.7%) (↑1.8%) (↑4.9%) (↑2.7%) (↑4.2%)
Table 2: Benchmark results for cross-dataset evaluation (Protocol‑1, frame-level AUC). All detectors are trained on FF++ c23(Rossler et al. 2019a) and evaluated on other deepfake datasets.
Methods Backbone UniFace BleFace MobSwap FaceDan InSwap SimSwap Avg.
F3Net (Qian et al. 2020) Xception 0.809 0.808 0.867 0.717 0.757 0.674 0.772
SPSL (Liu et al. 2021a) Xception 0.747 0.748 0.885 0.666 0.643 0.665 0.726
SRM (Luo et al. 2021) Xception 0.749 0.704 0.779 0.659 0.793 0.694 0.730
CORE (Ni et al. 2022) Xception 0.871 0.843 0.959 0.774 0.855 0.724 0.838
RECCE (Cao et al. 2022) Designed 0.898 0.832 0.925 0.848 0.848 0.768 0.853
SLADD (Chen et al. 2022) Xception 0.878 0.882 0.954 0.825 0.879 0.794 0.869
SBI (Shiohara and Yamasaki 2022) EfficientNet-B4 0.724 0.891 0.952 0.594 0.712 0.701 0.762
UCF (Yan et al. 2023) Xception 0.831 0.827 0.950 0.862 0.809 0.647 0.821
IID (Huang et al. 2023) Designed 0.839 0.789 0.888 0.844 0.789 0.644 0.799
LSDA (Yan et al. 2024a) EfficientNet-B4 0.872 0.875 0.930 0.721 0.855 0.793 0.841
ProDet (Cheng et al. 2024) EfficientNet-B4 0.908 0.929 0.975 0.747 0.837 0.844 0.873
CDFA (Lin et al. 2024b) SwinV2-B 0.762 0.756 0.823 0.803 0.772 0.757 0.779
ForensicsAdapter (Cui et al. 2024) CLIP (ViT-L/14) 0.969 0.886 0.963 0.943 0.937 0.917 0.936
Effort (Yan et al. 2024b) CLIP (ViT-L/14) 0.962 0.873 0.953 0.926 0.936 0.926 0.929
DPGNet (ours) CLIP (ViT-L/14) 0.987 0.984 0.990 0.974 0.972 0.984 0.982
(↑1.86%) (↑5.92%) (↑1.54%) (↑3.29%) (↑3.74%) (↑6.26%) (↑4.91%)
Table 3: Benchmarking Results for Cross-Method Evaluation (Protocol-2, Video-Level AUC). All detectors are trained on FF++ c23 (Rossler et al. 2019b) and evaluated on other deepfake datasets.
Evaluation Protocol.
The results in Figure 2 show that simply using high-confidence pseudo labels to align with the domain does not lead to improved performance, especially for detectors with lower performance, and may even lead to a decrease in performance. Accordingly, we adopt two widely used standard protocols for evaluation: Protocol 1 is used for cross-dataset evaluation, and Protocol 2 is used for cross-operation evaluation within the FF++ domain. For Protocol 1, the model is trained using the labeled source domain dataset FF++ and the small unlabeled dataset UCDDP, and the performance is evaluated on the test set corresponding to the UCDDP sampling dataset (DFDC, DFDCP, CDF-v1, CDF-v2, DFD) to evaluate the generalization ability across datasets. For Protocol 2, the model is trained using FF++ and the small unlabeled dataset UDF40, and the performance is evaluated on the test subset corresponding to the UDF40 sampling dataset DF40 to evaluate the generalization ability across different forgery methods under a consistent data distribution. To further demonstrate that our performance improvement is not due to the use of unlabeled data, we let the baseline methods use the same unlabeled dataset (UCDDP or UDF40) for additional comparison by generating and utilizing pseudo labels.
Implementation Details.
We adopt CLIP ViT-L/14 (Radford et al. 2021) as the visual backbone, with input images resized to $\times224$ pixels. During training, we sample 16 frames per video, while 32 frames are used for testing. The model is optimized using the Adam optimizer (Kingma 2014) with a learning rate of 0.00008 and a weight decay of 0.0005. For training, the batch size is set to 32 for the source domain (FF++) and 10 for unlabeled data (UCDDP or UD40), with a test batch size of 32. Standard data augmentation techniques, including random cropping and flipping, are applied to enhance data diversity. For feature library construction, we set the initial confidence threshold to $\lambda_{tf}=0.9$. The dynamic pseudo-labeling threshold $\lambda_{\text{lt}}$ starts at 0.85 and gradually decreases to 0.70 during training. Loss hyperparameters $\lambda$, $\lambda_{1}$, $\lambda_{2}$, and $\beta$ are empirically set to 0.8, 0.4, 0.5, and 0.1, respectively. For evaluation, we report frame-level and video-level Area Under the Curve (AUC), a standard metric in deepfake detection, to compare our method with prior work. AUC provides a robust measure of classification performance across varying thresholds. All experiments are conducted on an NVIDIA RTX 4090 GPU.
Methods Train Set $\lambda$ Cross-method Evaluation
UniFace BleFace MobSwap FaceDan InSwap SimSwap Avg.
F-Ada
FF++ - 0.919 0.818 0.940 0.904 0.904 0.856 0.890
+UDF40 0.9 0.924 0.837 0.946 0.912 0.916 0.876 0.902
+UDF40 0.8 0.929 0.837 0.947 0.917 0.919 0.862 0.902
+UDF40 0.7 0.924 0.835 0.947 0.912 0.921 0.867 0.901
Effort
FF++ - 0.940 0.825 0.911 0.883 0.907 0.885 0.892
+UDF40 0.9 0.932 0.852 0.918 0.897 0.899 0.889 0.898
+UDF40 0.8 0.938 0.837 0.928 0.896 0.908 0.899 0.901
+UDF40 0.7 0.940 0.841 0.932 0.897 0.910 0.901 0.904
Ours +UDF40 Adp 0.972 0.971 0.981 0.954 0.952 0.974 0.967
(↑3.2%) (↑11.9%) (↑4.9%) (↑3.7%) (↑4.2%) (↑7.3%) (↑6.3%)
Table 4: Cross-dataset evaluation of the baseline methods using unlabeled data (Frames-level AUC). U represents the unlabeled sampling data set corresponding to the test set, and $\lambda$ denotes the confidence threshold used in training.
Methods Train Set $\lambda$ Cross-dataset Evaluation
CDF-v1 CDF-v2 DFD DFDC DFDCP Avg.
F-Ada
FF++ - 0.914 0.900 0.933 0.843 0.890 0.896
+UCDDP 0.9 0.903 0.905 0.924 0.835 0.874 0.888
+UCDDP 0.8 0.936 0.906 0.927 0.834 0.868 0.894
+UCDDP 0.7 0.924 0.897 0.867 0.842 0.885 0.883
Effort
FF++ - 0.926 0.872 0.922 0.822 0.835 0.875
+UCDDP 0.9 0.924 0.907 0.929 0.833 0.845 0.888
+UCDDP 0.8 0.935 0.901 0.920 0.840 0.842 0.888
+UCDDP 0.7 0.933 0.891 0.912 0.829 0.830 0.879
Ours +UCDDP Adp 0.973 0.957 0.951 0.892 0.917 0.938
(↑3.7%) (↑5.0%) (↑2.2%) (↑5.2%) (↑2.7%) (↑4.2%)
Table 5: Cross-method evaluation of baseline methods using unlabeled data (Frames-level AUC). U represents the unlabeled sampling data set corresponding to the test set, and $\lambda$ denotes the confidence threshold used in training.
Number of samples CDF-v1 CDF-v2 DFDC DFDCP DFD
AUC AP EER AUC AP EER AUC AP EER AUC AP EER AUC AP EER
$6k$ 0.975 0.984 8.0 0.940 0.968 13.4 0.862 0.889 22.3 0.876 0.936 20.6 0.947 0.993 11.1
$12k$ 0.985 0.990 6.6 0.952 0.975 11.5 0.867 0.891 21.9 0.891 0.943 19.5 0.939 0.992 12.7
$18k$ 0.973 0.985 8.7 0.957 0.978 11.0 0.892 0.914 19.1 0.917 0.956 16.8 0.951 0.994 10.4
$24k$ 0.975 0.984 8.6 0.964 0.979 9.6 0.883 0.906 20.0 0.910 0.952 17.0 0.946 0.993 11.1
Table 6: Ablation study on the number of samples of unlabeled datasets, evaluated using frame-level AUC, AP, and EER.
Number of samples UniFace BleFace MobSwap FaceDan InSwap SimSwap
AUC AP EER AUC AP EER AUC AP EER AUC AP EER AUC AP EER AUC AP EER
$4k$ 0.959 0.965 8.4 0.938 0.944 12.1 0.973 0.995 7.0 0.928 0.934 12.9 0.947 0.942 10.2 0.955 0.960 9.1
$8k$ 0.970 0.971 7.3 0.964 0.963 8.7 0.976 0.995 6.9 0.938 0.940 11.3 0.949 0.942 10.3 0.968 0.967 7.6
$12k$ 0.972 0.971 7.7 0.968 0.970 8.2 0.978 0.996 6.6 0.954 0.958 10.1 0.953 0.950 10.1 0.973 0.975 6.9
$16k$ 0.973 0.973 7.9 0.971 0.972 7.3 0.981 0.997 6.4 0.963 0.964 8.8 0.957 0.954 9.1 0.976 0.994 6.5
Table 7: Ablation study on the number of samples of unlabeled datasets, evaluated using frame-level AUC, AP, and EER.
Ours Cross Cross Avg.
TCA CPG CD -dataset -method
$\times$ $\times$ $\times$ 0.871 0.857 0.864
$\times$ 0.896 0.927 0.912
$\times$ $\times$ 0.903 0.924 0.914
$\times$ $\times$ 0.926 0.950 0.938
$\times$ 0.924 0.956 0.940
$\times$ 0.934 0.958 0.946
0.938 0.967 0.953
Table 8: Ablation study on core components. Results for Cross-dataset (UCDDP) and Cross-method (UDF40).
Ours Cross Cross Avg.
$\mathcal{L}_{\rm alig}$ $\mathcal{L}_{\rm con}$ $\mathcal{L}_{\rm dis}$ -dataset -method
$\times$ $\times$ $\times$ 0.876 0.857 0.867
$\times$ $\times$ 0.919 0.946 0.933
$\times$ 0.924 0.946 0.935
$\times$ 0.934 0.952 0.943
0.938 0.967 0.953
Table 9: Ablation study on embedding alignment ($\mathcal{L}{\rm alig}$), contrast enhancement ($\mathcal{L}{\rm con}$), and cross-domain distillation ($\mathcal{L}_{\rm dis}$). Results for cross-dataset (UCDDP) and cross-method (UDF40).
Methods Cross-dataset Cross-method Avg.
Ori 0.872 0.857 0.865
DANN 0.846 0.838 0.842
NAMC 0.878 0.877 0.878
SDAT 0.864 0.842 0.853
Ours 0.938 0.967 0.953
Table 10: Comparison with domain adaptation methods.
Detection Performance
Table 2 presents the results of cross-dataset evaluation under Protocol-1. DPGNet achieves an average frame-level AUC of 0.938, surpassing the best baseline, ForensicsAdapter (0.896), by 4.2%. Notably, DPGNet excels on challenging datasets such as DFDC (AUC of 0.892, +4.9% over ForensicsAdapter) and CDF-v2 (AUC of 0.957, +5.7%). This superior performance is attributed to DPGNet’s text-guided alignment technique, which leverages text embeddings to capture domain-invariant forgery cues and effectively utilizes unlabeled data via a curriculum-driven pseudo labele generation technique. For Protocol-2, Table 3 reports cross-method evaluation results on DF40, where DPGNet achieves an average video-level AUC of 0.982, outperforming ForensicsAdapter (0.936) by 4.91%. DPGNet demonstrates significant improvements on advanced forgery methods, such as SimSwap (AUC of 0.984, +6.26%) and BleFace (AUC of 0.984, +5.92%), highlighting its robustness to diverse manipulation techniques within a consistent domain.
To assess the impact of leveraging unlabeled data, we compare DPGNet against baselines augmented with pseudo-labeling at fixed confidence thresholds (Tables 4 and 5). For cross-dataset evaluation, ForensicsAdapter with UCDDP (0.7 threshold) suffers a performance drop (average AUC of 0.883, -1.3% compared to FF++ alone), likely due to noisy pseudo labels. In contrast, DPGNet with UCDDP achieves a robust AUC of 0.938, demonstrating its ability to prioritize high-value samples through dynamic curriculum learning. Similarly, in cross-method evaluation, DPGNet with UDF40 achieves an average AUC of 0.967, surpassing ForensicsAdapter (0.902, +6.5%) and Effort (0.904, +6.3%). We may notice that the baseline performance of UCDDP and UDF40 has limited improvement and minimal variation across fixed thresholds, which stems from the limited sample size of these datasets. After threshold-based filtering, the number of usable training samples is further reduced, limiting the impact on models initialized with pre-trained weights, which are inherently stable to small data increments.
Figure 4: Cross-dataset (left) and cross-method evaluation.
Ablation Study
To dissect the contributions of the DPGNet design, we performed ablation studies on the unlabeled sample size, key components, loss functions, and unsupervised baselines, providing insights into its robust generalization capabilities across domains and forgery methods.
Unlabeled Sample Size. Tables 4 and 5 evaluate the effect of varying unlabeled sample sizes from UCDDP and UDF40. In cross-dataset evaluation, increasing UCDDP samples from 6k to 24k improves the average AUC from 0.896 to 0.938, stabilizing at 18k samples (e.g., DFDCP: AUC 0.914, EER 16.9). In cross-method evaluation, scaling UDF40 samples from 4k to 16k raises the average AUC from 0.950 to 0.970 (e.g., MobSwap: AUC 0.981, EER 6.4). These results demonstrate that our method achieves significant performance gains in target domains with minimal unlabeled data, particularly in cross-method detection, where small sample sizes yield substantial improvements, highlighting the sample efficiency of our domain adaptation approach for deep face forgery detection.
Core Components. Table 8 evaluates the core components of DPGNet: text-guided cross-domain alignment (TCA) and curriculum-driven pseudo label generation (CPG). Additionally, we provide a more in-depth evaluation of the cross-domain distillation (CD) strategy. The baseline without these components yields an AUC of 0.867. TGA alone boosts performance to 0.938 (+7.1%) by aligning visual-textual embeddings, mitigating domain gaps. Adding CPL increases the AUC to 0.946 (+7.9%) by progressively lowering the pseudo-labeling threshold to include challenging samples, while CD ensures robustness, maintaining the AUC at 0.953. This synergy drives its superior generalization.
Ablation on Loss Functions. Table 9 evaluates the loss components of DPGNet: embedding alignment ($\mathcal{L}_{\text{alig}}$), contrastive enhancement ($\mathcal{L}_{\text{con}}$) and distillation across domain ($\mathcal{L}_{\text{dis}}$). The baseline without these losses achieves an AUC of 0.867. Adding $\mathcal{L}_{\text{alig}}$ improves the AUC to 0.933 (+6.6%) by ensuring domain-invariant representations. Including $\mathcal{L}_{\text{con}}$ and $\mathcal{L}_{\text{dis}}$ further stabilizes the performance to 0.953.
Comparison with Unsupervised Methods. Table 10 compares DPGNet against unsupervised methods: DANN (Ganin and Lempitsky 2015), NAMC(Zhou et al. 2024), and source-free domain adaptation(SDAT). DPGNet outperforms these methods, utilizing text-guided alignment and curriculum learning to capture diverse forgery patterns while preserving source domain knowledge.
Figure 5: T-SNE visualization on the cross-method test.
Feature Distribution Visualization
To show the uniqueness of DPGNet from the feature distribution level, we use T-SNE (Van der Maaten and Hinton 2008) to visualize the feature distribution of the baseline and DPGNet. As shown in Figure 5, the baseline model exhibits significant overlap between real and fake features, indicating that it lacks semantic distinction in the target domain. In contrast, DPGNet learns more robust true/false semantics and significantly increases the separation between features of different categories. This larger separation surface DPGNet can more effectively bridge the gap between the source and target domains, thereby improving performance.
Conclusion
This work addresses the critical challenge of detecting deepfakes in realistic settings, where vast amounts of unlabeled data remain underutilized. We propose DPGNet, a novel framework that leverages text-guided alignment, curriculum-driven pseudo label generation, to fully exploit unlabeled deepfakes. By unifying visual and textual embeddings in a domain-invariant space, DPGNet captures generalizable features, dynamically selects informative samples to avoid overfitting, and preserves source-domain robustness via distillation. Extensive benchmarks (11 popular datasets) show that DPGNet consistently outperforms state-of-the-art methods (+6.3%). In the future, we plan to extend DPGNet with incremental learning and explore its applicability to related tasks such as anomaly detection and face anti-spoofing.